OpenStudy (anonymous):

what is the slope intercept form for points 6,8 10,3 please explain

OpenStudy (anonymous):

@texaschic101

OpenStudy (anonymous):

First you have to find the slope m using (y2-y1)/(x2-x1), where 6 is x1, 8 is y1, 10 is x2, and 3 is y2

OpenStudy (anonymous):

please i need helpp

OpenStudy (jdoe0001):

http://mathb.in/12756 <--- what would be the value of the slope?

OpenStudy (anonymous):

slope = (3-8)/(10-6). what is the slope?

OpenStudy (anonymous):

-5/4 @larryboxaplenty

OpenStudy (anonymous):

Good, now we put in into slope intercept form, We can choose either points, I will pick (6,8). y-y1=m(x-x1) where y1 is 8, m is (-5/4), and x1 is 6

OpenStudy (anonymous):

Plug it x1 and y1, and solve for y.

OpenStudy (anonymous):

since -5/4 =-1.25 i out that in as m

OpenStudy (anonymous):

@larryboxaplenty

OpenStudy (anonymous):

y-(8)=(-5/4)(x-6) solve for y

OpenStudy (anonymous):

can u walk me throught the steps of that

OpenStudy (anonymous):

y-8=(-5/4)(x-6) keep y-8 alone for now, and distribute the (-5/4) to the x and -6

OpenStudy (anonymous):

What do you get

OpenStudy (anonymous):

for -5/4 x6 i got 7.-5 and the x just stays right

OpenStudy (anonymous):

Kind of hard to read, but yes. You get (-5/4)x+7.5

OpenStudy (anonymous):

Now you have y-8=(-5/4)x+7.5 You want the y by itself on the left side so solve for y

OpenStudy (anonymous):

i really dont know what to do from here

OpenStudy (anonymous):

You want to get rid of the negative 8 on the left side, so you add 8 on both sides of the equation.

OpenStudy (anonymous):

visual please. i am a visual learner

OpenStudy (anonymous):

y-8+8=(-5/4)x+7.5+8 -8+8=0, so: y=(-5/4)x+7.5+8 y=(-5/4)x+15.5 that is the equation.

OpenStudy (anonymous):

do you understand?

OpenStudy (anonymous):

yes now i do thanks you

OpenStudy (anonymous):

Remember, if you have 2 points: 1)find slope 2)use a point in slope-intercept form 3)solve for y